1.6. Write in both polar and rectangular form: (a) eln(5); (b) in eptic
Answer #1

9 that a logo bloga ens e loge - s loge=5 i ens; = si annarris) 1 Rectangular form - a tib. here, a=0, b=s. .: Otsi i.e, Singd leatio) let, 2= d ledtid) de deQ(1+i) do C allti) 2= (1+i).et we know that, eio=coso toisino. :2= (1+i) eatai I = eq eli ti